Compute the first three derivatives of f(x)=2x^5+x^(3/5)-1/2x
1
Expert's answer
2019-05-13T12:22:58-0400
"f(x)=2x^5+x^{3/5}-1/2\cdot x^{-1}"

First derivative:

"f'(x)=10x^4+3/5\cdot x^{-2/5}+1/2\cdot x^{-2}"

Second derivative:

"f''(x)=40x^3-6/25\cdot x^{-7/5}-x^{-3}"

Third derivative:

"f'''(x)=120x^2+42/125\cdot x^{-12/5}+3x^{-4}"